OpenStudy (surryyy):

Select the equation that contains the points, (-2, -2) and (4, 10). y-6=2(x-2) y-6= -2(x-2) y+6 = -2(x+1) y+2=6(x+6)

OpenStudy (texaschic101):

first we find the slope by using the slope formula. slope(m) = (y2 - y1) / (x2 - x1) (-2,-2)...x1 = -2 and y1 = -2 (4,10)....x2 = 4 and y2 = 10 now we sub slope(m) = (10 - (-2) / (4 - (-2) -- pay close attention to your signs slope(m) = (10 + 2) / (4 + 2) slope(m) = 12/6 which equals 2 Now we use point slope formula y - y1 = m(x - x1) slope(m) = 2 you can use either of your points... (-2,-2)...x = -2 and y = -2 now we sub y - (-2) = 2(x - (-2) y + 2 = 2(x + 2) <==== here is one possible answer using the other point (4,10)...x1 = 4 and y1 = 10 y - 10 = 2(x - 4) < ==== here is another possible answer choice hmmm...are you sure you have the right points ? Well....the only one that has a slope of 2 is A....it cannot be the others...unless your points are wrong.

OpenStudy (surryyy):

so a would be my answer?